# Data Fusion for Heterogeneous Treatment Effect Estimation with Multi-Task Gaussian Processes

### v1

@@ 3.2 Multi-task Gaussian Processes
In our setting, we will estimate $\tau(x)$ with a T-learner approach (Künzel et al., 2019), i.e. we individually estimate $\mathbb{E}[Y^{0}\mid\mathbf{x}]$ and $\mathbb{E}[Y^{1}\mid\mathbf{x}]$ with separate models. For brevity, we will focus on $\mathbb{E}[Y^{1}\mid\mathbf{x}]$ in the exposition. We then write the study specific response surfaces as:
@@
For $\mathbf{f}(\mathbf{x})=(f^{\text{e}}\left(\mathbf{x}\right),f^{\text{o}}\left(\mathbf{x}\right))$, we assign a multi-task GP prior $\textbf{f}\sim\mathcal{GP}\left(\boldsymbol{\mu},K\right)$, where $\boldsymbol{\mu}$ is a bivariate mean function (which we again take to be 0) and $K$ is a positive $2\times 2$ matrix-valued covariance function that maps input points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ to matrices quantifying the relationship of the outputs at these inputs. We will introduce a specific choice of $K$ in the next section. They key strength of multi-task GP s is the ability to share information between tasks, which we will leverage for data fusion. We provide an illustrative comparison to independent GP s in Appendix A.2.
As usual, we assume a Gaussian observation model:
$y_{i}^{\text{e}}=f^{\text{e}}\left(\mathbf{x}_{i}^{\text{e}}\right)+\epsilon_{i}^{\text{e}},\quad y_{j}^{\text{o}}=f^{\text{o}}\left(\mathbf{x}_{j}^{\text{o}}\right)+\epsilon_{j}^{\text{o}}$
where $\epsilon_{i}^{\text{e}}\sim\mathcal{N}(0,\sigma^{2})$ and $\epsilon_{j}^{\text{o}}\sim\mathcal{N}(0,\sigma^{2})$ independently across tasks and observations for $i=1,\ldots,n^{\text{e}}$ and $j=1,\ldots,n^{\text{o}}$. For simplicity, we have assumed a common variance $\sigma^{2}$ between tasks but this can be easily extended to task-specific variances. We write $\mathbf{y}^{\text{e}}$ for the $n^{\text{e}}$-vector of outcomes and $X^{\text{e}}$ for the $(n^{\text{e}}\times p)$ matrix of covariates respectively, and similarly for the observational dataset.
In data fusion settings for clinical trials, it is usually the case that $n^{\text{e}}\ll n^{\text{o}}$, and $p(\mathbf{x}^{\text{e}})$ will have a smaller support than $p(\mathbf{x}^{\text{o}})$. Although the RCT allows for unbiased estimation, it may not be representative of the population of interest. Making inferences in the target population thus usually requires some level of extrapolation, which relies on correct model specification and other strong untestable assumptions. Our goal is thus to leverage the multi-task GP to carefully extrapolate $f^{\text{e}}$ to $\mathbf{x}$ values outside the support of $p(\mathbf{x}^{\text{e}})$, given the observations $(\mathbf{y}^{\text{o}},X^{\text{o}})$. By sharing information between the tasks and leveraging the flexible nature of GP s, we hope to increase precision of our estimates without introducing bias. The object of interest is then the posterior distribution of $f^{\text{e}}$ given $\mathcal{D}=(\mathbf{y}^{\text{e}},X^{\text{e}},\mathbf{y}^{\text{o}},X^{\text{o}})$, which is also a GP from conjugacy. Crucially, the Bayesian framework will enable reliable uncertainty estimates when extrapolating.
@@ 3.3 Intrinsic coregionalisation models (ICMs)
The ICM is a special case of a multi-task GP with a separable kernel, which is particularly interpretable and suitable for data fusion (Vargas-Guzman and Warrick, 1999). Here, the functions are linear combinations of independent random functions, which implies a particularly simple covariance function. For our setting, we have
$f^{\text{e}}\left(\mathbf{x}\right)=\sum_{q=1}^{Q}\alpha_{q}^{\text{e}}u_{q}\left(\mathbf{x}\right),\quad f^{\text{o}}\left(\mathbf{x}\right)=\sum_{q=1}^{Q}\alpha_{q}^{\text{o}}u_{q}\left(\mathbf{x}\right)$
@@
One can easily show that the multi-task covariance function can be expressed as
@@
$B$ $=\begin{bmatrix}\beta^{\text{e}}&\beta^{\text{eo}}\\ \beta^{\text{eo}}&\beta^{\text{o}}\end{bmatrix}=\begin{bmatrix}\left(\alpha_{1}^{\text{e}}\right)^{2}+\left(\alpha_{2}^{\text{e}}\right)^{2}&\alpha_{1}^{\text{e}}\alpha_{1}^{\text{o}}+\alpha_{2}^{\text{e}}\alpha_{2}^{\text{o}}\\[1.42262pt] \alpha_{1}^{\text{e}}\alpha_{1}^{\text{o}}+\alpha_{2}^{\text{e}}\alpha_{2}^{\text{o}}&\left(\alpha_{1}^{\text{o}}\right)^{2}+\left(\alpha_{2}^{\text{o}}\right)^{2}\end{bmatrix}$ (1)
The posterior distribution over $f^{\text{e}}$ can now be computed using standard theory (see Appendix A.3). For a given test point $\mathbf{x}_{*}$, we have that the posterior distribution is Gaussian
@@
where $k(\mathbf{x}_{*},X^{s})$ is the row vector of covariances between $\mathbf{x}_{*}$ and $X^{s}$, and $k^{\text{e}}\left(X,\mathbf{x}_{*}\right)=k^{\text{e}}(\mathbf{x}_{*},X)^{T}$. Given this posterior, we can report $m^{\text{e}}(\mathbf{x}_{*})$ as the point estimate of the response surface at $\mathbf{x}_{*}$, with associated posterior variance $V^{\text{e}}(\mathbf{x}_{*})$ which can be used to compute credible intervals.
@@ 3.4 Causal-ICM
[…]
@@
Given the above parameterization, $\rho$ can be interpreted as the maximum degree of borrowing allowed, where higher values of $|\rho|$ indicate a higher degree of borrowing. Note that while the above setting implies $\beta^{\text{e}}=\beta^{\text{o}}=1$ (and $\beta_{\text{eo}}=\rho$), this will not impede the flexibility of the model much as long as the kernel $k(\mathbf{x},\mathbf{x})$ includes an independent scaling term (such as the variance for an RBF kernel) and the scales of $f^{\text{e}}$ and $f^{\text{o}}$ are not too different.
Setting $\alpha_{2}^{\text{e}}=0$ allows us to easily set the other ICM coefficients to interpretable values given the desired value of $\rho$. With our suggested settings of the ICM coefficients, we then have:
@@
We now have that $f^{\text{e}}=u_{1}$ and $f^{\text{o}}$ is a scaled version of $f^{\text{e}}$ plus a term which essentially encompasses the confounding effect. Choosing a positive $\rho\in(0,1)$ is reasonable in this case, as we expect the sign of $f^{\text{o}}(\mathbf{x})$ to be highly correlated with $f^{\text{e}}(\mathbf{x})$; both tasks model the same quantity - the conditional expectation of the outcome - with the exception that $f^{o}$ is contaminated with confounding. We find that this restriction of $\alpha_{2}^{\text{e}}$ does not affect flexibility too much in practice, whilst greatly improving the simplicity and interpretability of the coefficients. The remaining hyperparameters, such as the kernel hyperparameters and sampling variance, are then estimated via marginal likelihood maximisation.
@@ 3.4.1 Variance bound for conditional mean function under Causal-ICM
When evaluating treatment effects to inform decision making, it is crucial to provide a measure of uncertainty alongside point estimates. For example, when comparing the effectiveness of a drug against a placebo, a regulator may insist that the treatment effect must be significantly positive before providing approval. In statistical terms, the credible interval (for some significance level) must not contain zero. The multi-task Gaussian process framework provides a natural mechanism by which to quantify uncertainty via the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$.
For data fusion, a valid concern is that the information in the observational data ‘swamps’ that of the experimental data. More precisely, there is a danger that as the sample size of the observational study increases, estimates for the true treatment effect $\tau(\mathbf{x})$ become increasingly biased, and more problematically, increasingly confident around this biased estimate. We now demonstrate that the ICM is in fact robust against this behaviour, via a lower bound for the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$ as the number of observational data points $n^{\text{o}}$ increases.
@@
where $\rho\in(0,1)$ is the specified borrowing hyperparameter.
Proof.
Let $n^{\textnormal{e}},n^{\textnormal{o}}\geq 0$ denote the sizes of the experimental and observational datasets respectively. We outline the proof for $n^{\textrm{e}}=0$ here, which corresponds to the prior-only case for the RCT. The proof for the general case $n^{\textrm{e}}>0$ is deferred to Appendix A.4 but uses a similar technique. The key is that when $n^{\textrm{e}}=0$, it is possible to write the posterior variance as
@@
where $V^{\textnormal{e}}_{\mathcal{D}_{\textnormal{o}}}$ the posterior variance for a GP with kernel $k(\mathbf{x},\mathbf{x})$ fit only to $\mathcal{D}_{\textnormal{o}}=(\mathbf{y}^{\textnormal{o}},X^{\textnormal{o}})$, and $k(\mathbf{x}_{*},\mathbf{x}_{*})$ is the prior variance. As $V^{\textnormal{e}}_{\mathcal{D}_{\textnormal{o}}}\left(\mathbf{x}_{*}\right)\geq 0$, we obtain the desired lower bound. ∎
We now provide the key interpretation of proposition above: even as $n^{\text{o}}\to\infty$, the posterior variance of $f^{\text{e}}(\mathbf{x}_{*})$ conditional on the full dataset will at most decrease by a factor of $(1-\rho^{2})$ relative to the posterior variance given the experimental data only, where $|\rho|\leq 1$. We conjecture that this inequality is tight as $n^{\text{o}}\to\infty$. This can also be interpreted as limiting the ‘effective sample size’ of the observational dataset to be $1/\left(1-\rho^{2}\right)$ relative to the experimental dataset.
@@ 3.4.2 Tuning $\rho$
We now outline a data-adaptive procedure to select $\rho$, where the goal is to prevent information from the large confounded observational study from swamping that of the smaller unconfounded RCT. Our proposal is based on cross-validation to minimize the Root Mean Squared Error (RMSE) on weighted held-out RCT patients only, where the weighting tailors $\rho$ for extrapolation. Specifically, we propose to use 5-fold cross-validation, as the RCT sample size is usually small, and we consider $\rho\in[0.0,0.1,\cdots,1.0]$ on a grid. In particular, we minimize the following objective:
@@
where $\{\tilde{y}_{i},\tilde{\mathbf{x}}_{i}\}_{i=1}^{n_{\text{held-out}}}$ is a held-out subset of $(\mathbf{y}^{\textnormal{e}},X^{\textnormal{e}})$ that was not used to train $m^{\text{e}}$. The above objective is motivated as follows. Firstly, we only evaluate predictions on RCT patients where there is no confounding present. Secondly, the weights will upweight RCT patients who have higher probability of arising from the observational dataset, so $\rho$ will be tailored to improve extrapolation beyond the RCT support. We will see in the experiments that this works well in practice.
@@ 3.4.3 Extension to CATE Estimation
When we have both treatment and control arms, in order to focus on the sharing between the experimental and observational datasets, we use two separate rank-2 ICMs to estimate $(f_{0}^{\text{e}},f_{0}^{\text{o}})$ and $(f_{1}^{\text{e}},f_{1}^{\text{o}})$, giving us independent posteriors for $f_{0}^{\text{e}}$ and $f_{1}^{\text{e}}$ which we combine to obtain the posterior on $\tau(\mathbf{x})$. As the posteriors are independent GP s, we can simply compute the difference of the means and sum the variances appropriately. We also allow separate kernel hyperparameters between the two ICM s, but select a common $\rho$ between the two arms by choosing the value that minimizes the average of $\mathcal{L}\left(\rho\right)$ across the two treatment groups. As mentioned earlier, this is essentially a T-learner approach (Künzel et al., 2019). In practice, one could also investigate the usage of a rank-4 ICM to allow for cross-pooling, but our choice is computationally expedient and performs well.

### v2

@@ 3.2 Multi-task Gaussian Processes
In our setting, we use a T-learner approach (Künzel et al., 2019) to estimate $\tau(\mathbf{x})$, i.e. we individually estimate $\mathbb{E}[Y^{0}\mid\mathbf{x}]$ and $\mathbb{E}[Y^{1}\mid\mathbf{x}]$ with separate models. For brevity, we will focus on $\mathbb{E}[Y^{1}\mid\mathbf{x}]$ in the exposition. We write the study specific response surfaces as:
@@
For $\mathbf{f}(\mathbf{x})=(f^{\text{e}}\left(\mathbf{x}\right),f^{\text{o}}\left(\mathbf{x}\right))$, we assign a multi-task GP prior $\textbf{f}\sim\mathcal{GP}\left(\boldsymbol{\mu},K\right)$, where $\boldsymbol{\mu}$ is a bivariate mean function (taken again to be 0) and $K$ is a positive $2\times 2$ matrix-valued covariance function that maps input points $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ to matrices quantifying the relationship of the outputs at these inputs. We specify $K$ in the next section. The key strength of multi-task GP s is the ability to share information between tasks, which we will leverage for data fusion. See Appendix B Fig 3 for an illustrative comparison to independent GP s.
We assume a Gaussian observation model,
$y_{i}^{\text{e}}=f^{\text{e}}\left(\mathbf{x}_{i}^{\text{e}}\right)+\epsilon_{i}^{\text{e}},\quad y_{j}^{\text{o}}=f^{\text{o}}\left(\mathbf{x}_{j}^{\text{o}}\right)+\epsilon_{j}^{\text{o}},$
where $\epsilon_{i}^{\text{e}},\epsilon_{j}^{\text{o}}\sim\mathcal{N}(0,\sigma^{2})$ independently across tasks and observations for $i=1,\ldots,n^{\text{e}}$ and $j=1,\ldots,n^{\text{o}}$. For simplicity, we assume a common variance $\sigma^{2}$ between tasks but this can be easily extended to task-specific variances. We write $\mathbf{y}^{\text{e}}$ for the $n^{\text{e}}$-vector of outcomes and $X^{\text{e}}$ for the $(n^{\text{e}}\times p)$ matrix of covariates respectively, and similarly for the observational dataset.
In data fusion settings for clinical trials, typically $n^{\text{e}}\ll n^{\text{o}}$, and $p(\mathbf{x}^{\text{e}})$ may not have support over the entire target population. Making inferences across the whole population thus requires some level of extrapolation. We leverage the multi-task GP to improve estimation of $f^{\text{e}}$ outside the support of $p(\mathbf{x}^{\text{e}})$, given the observations $(\mathbf{y}^{\text{o}},X^{\text{o}})$. By sharing information between tasks and leveraging the flexible nature of GPs, we hope to increase precision of our estimates without introducing significant bias. The object of interest is the posterior distribution of $f^{\text{e}}$ given $\mathcal{D}=(\mathbf{y}^{\text{e}},X^{\text{e}},\mathbf{y}^{\text{o}},X^{\text{o}})$, which, by conjugacy, is also a GP. Crucially, the Bayesian framework also provides reliable uncertainty estimates in regions of extrapolation.
@@ 3.3 Intrinsic coregionalisation models (ICMs)
The Intrinsic Coregionalization Model (ICM) is a special case of a multi-task GP with a separable kernel, which is particularly interpretable and suitable for data fusion (Vargas-Guzman and Warrick, 1999). Each task is a linear combination of independent latent functions, yielding a particularly simple covariance structure. For our setting, we have
$f^{\text{e}}\left(\mathbf{x}\right)=\sum_{q=1}^{Q}\alpha_{q}^{\text{e}}u_{q}\left(\mathbf{x}\right),\quad f^{\text{o}}\left(\mathbf{x}\right)=\sum_{q=1}^{Q}\alpha_{q}^{\text{o}}u_{q}\left(\mathbf{x}\right),$
@@
The multi-task covariance function can be expressed as
@@
$B$ $=\begin{bmatrix}\beta^{\text{e}}&\beta^{\text{eo}}\\ \beta^{\text{eo}}&\beta^{\text{o}}\end{bmatrix}=\begin{bmatrix}\left(\alpha_{1}^{\text{e}}\right)^{2}+\left(\alpha_{2}^{\text{e}}\right)^{2}&\alpha_{1}^{\text{e}}\alpha_{1}^{\text{o}}+\alpha_{2}^{\text{e}}\alpha_{2}^{\text{o}}\\[1.42262pt] \alpha_{1}^{\text{e}}\alpha_{1}^{\text{o}}+\alpha_{2}^{\text{e}}\alpha_{2}^{\text{o}}&\left(\alpha_{1}^{\text{o}}\right)^{2}+\left(\alpha_{2}^{\text{o}}\right)^{2}\end{bmatrix}$. (1)
The posterior distribution over $f^{\text{e}}$ can then be computed using standard theory (see Appendix C). For a given test point $\mathbf{x}_{*}$, we have that the posterior distribution is Gaussian
@@
where $k(\mathbf{x}_{*},X^{s})$ is the row vector of covariances between $\mathbf{x}_{*}$ and $X^{s}$, and $k^{\text{e}}\left(X,\mathbf{x}_{*}\right)=k^{\text{e}}(\mathbf{x}_{*},X)^{T}$. Given this posterior, $m^{\text{e}}(\mathbf{x}_{*})$ provides point estimates of the response surface, with associated posterior variance $V^{\text{e}}(\mathbf{x}_{*})$ which can be used to compute credible intervals.
@@ 3.4 Causal-ICM
We now introduce Causal-ICM. Specifically, we propose an interpretable parametrisation of the ICM for causal inference and introduce a bespoke procedure for learning its coefficients, rather than relying on standard marginal likelihood maximisation. This serves to control the influence of the observational dataset and avoid “swamping” the RCT data. Causal-ICM consists of the following choices for the ICM coefficients:
@@
Under this parameterization, $\rho\in(0,1)$ controls the degree of borrowing, with $\rho\to 1$ corresponding to maximum borrowing and $\rho\to 0$ to none (We provide an additional interpretation of $\rho$ in Appendix D, Proposition D.1). Although this implies $\beta^{\text{e}}=\beta^{\text{o}}=1$ (and $\beta_{\text{eo}}=\rho$), the model remains flexible as long as the kernel $k(\mathbf{x},\mathbf{x})$ includes an independent scaling term (e.g., the variance in an RBF kernel) and the scales of $f^{\text{e}}$ and $f^{\text{o}}$ are similar.
With these suggested settings for the ICM coefficients, we then have:
@@
Thus, $f^{\text{e}}=u_{1}$ and $f^{\text{o}}$ is a scaled version of $f^{\text{e}}$ plus a term capturing confounding. Choosing a positive $\rho\in(0,1)$ is reasonable, as $f^{\text{o}}(\mathbf{x})$ is expected to be highly correlated with $f^{\text{e}}(\mathbf{x})$ - both model the conditional expectation of the outcome - with $f^{\text{o}}$ additionally contaminated by confounding. In practice, restricting $\alpha_{2}^{\text{e}}$ minimally affects flexibility while improving interpretability. Remaining hyperparameters, including kernel parameters and sampling variance, are estimated via marginal likelihood maximisation.
@@ 3.4.1 Variance bound for conditional mean function under Causal-ICM
When evaluating treatment effects, it is important to quantify uncertainty alongside point estimates. The multi-task Gaussian process naturally provides this uncertainty via the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$. In data fusion, a key concern is that large observational samples may “swamp” the experimental data, yielding biased and overconfident estimates of $\tau(\mathbf{x})$. Causal-ICM mitigates this, as shown by a lower bound on the posterior variance $V^{\text{e}}(\mathbf{x}_{*})$ even as $n^{\text{o}}$ grows. This variance bound can also be seen as formally limiting the information flow from the observational dataset, similar to cut models (Lin et al., 2025a).
@@
where $\rho\in(0,1)$ is the borrowing hyperparameter.
Proof 3.2.
Let $n^{\textnormal{e}},n^{\textnormal{o}}\geq 0$ denote the sizes of the experimental and observational datasets respectively. We outline the proof for $n^{\textrm{e}}=0$ here, which corresponds to the prior-only case for the RCT. The proof for the general case $n^{\textrm{e}}>0$ is deferred to Appendix E. The key is that when $n^{\textrm{e}}=0$, the posterior variance can be written as
@@
where $V^{\textnormal{e}}_{\mathcal{D}_{\textnormal{o}}}$ is the posterior variance for a GP with kernel $k(\mathbf{x},\mathbf{x})$ fit only to $\mathcal{D}_{\textnormal{o}}=(\mathbf{y}^{\textnormal{o}},X^{\textnormal{o}})$, and $k(\mathbf{x}_{*},\mathbf{x}_{*})$ is the prior variance. As $V^{\textnormal{e}}_{\mathcal{D}_{\textnormal{o}}}\left(\mathbf{x}_{*}\right)\geq 0$, we obtain the desired lower bound.
An intuitive understanding of the result is as follows: for two random variables with correlation $\rho$, knowledge of one random variable does not inform us of the exact value of the other unless $\rho=1$. The key interpretation of Proposition 3.1 in the data fusion context is as follows: even as $n^{\text{o}}\to\infty$, the posterior variance of $f^{\text{e}}(\mathbf{x}_{*})$ conditional on the full dataset will at most decrease by a factor of $(1-\rho^{2})$ relative to the posterior variance given the experimental data only, where $|\rho|\leq 1$. We conjecture that this inequality is tight as $n^{\text{o}}\to\infty$. This can be interpreted as limiting the ‘effective sample size’ of the observational dataset to be $1/\left(1-\rho^{2}\right)$ relative to the experimental dataset. Crucially, this variance bound protects from having very tight credible intervals centred around a biased estimate - the posterior uncertainty will accurately reflect our distrust of the observational dataset. We also expect this ‘effective sample size’ effect to hold for the posterior mean of $f^{\text{e}}(\mathbf{x}_{*})$, where there the bias introduced to the posterior mean by $\mathcal{D}_{\text{o}}$ is limited again by the choice of $\rho$. We leave this interesting direction for future work. Although our focus is on the conditional mean functions, our approach also provides the posterior distribution of the confounding function, offering additional insights that may be of independent interest (see Appendix F).
@@ 3.4.2 Tuning $\rho$
We now outline a data-adaptive procedure to select $\rho$, where the goal is to prevent information from the large confounded observational study from swamping that of the smaller unconfounded RCT. Our proposal is based on cross-validation to minimize the Root Mean Squared Error (RMSE) on weighted held-out RCT patients only, where the weighting tailors $\rho$ for extrapolation. Specifically, use 5-fold cross-validation, as the RCT sample size is usually small, and we consider $\rho\in[0.0,0.1,\cdots,1.0]$ on a grid. In particular, we minimize the following objective:
@@
where $\{\tilde{y}_{i},\tilde{\mathbf{x}}_{i}\}_{i=1}^{n_{\text{held-out}}}$ is a held-out subset of $(\mathbf{y}^{\textnormal{e}},X^{\textnormal{e}})$ that was not used to train $m^{\text{e}}$. The above objective is motivated as follows. Firstly, we only evaluate predictions on RCT patients where there is no confounding present. Secondly, the weights will upweight RCT patients who have higher probability of arising from the observational dataset, so $\rho$ will be tailored to improve extrapolation beyond the RCT support. We then expect the largest gains in power of $\rho$ is chosen close to 1, indicating that the confounding effect is small. We will see in the experiments that this works well in practice.
@@ 3.4.3 Extension to CATE Estimation
When both treatment and control arms are available, we fit two separate rank-2 ICMs to model $(f_{0}^{\text{e}},f_{0}^{\text{o}})$ and $(f_{1}^{\text{e}},f_{1}^{\text{o}})$, focusing on sharing information between the experimental and observational datasets. This yields independent posteriors for $f_{0}^{\text{e}}$ and $f_{1}^{\text{e}}$, which we combine to obtain the posterior of $\tau(\mathbf{x})$ by differencing the means and summing the variances. We allow distinct kernel hyperparameters for the two ICMs while using a common $\rho$, chosen to minimise the average $\mathcal{L}(\rho)$ across treatment groups. This setup corresponds to a T-learner approach (Künzel et al., 2019).